OpenStudy (ksaimouli):

find the domain in f(x)=x2+4

OpenStudy (anonymous):

All real numbers

OpenStudy (ksaimouli):

show me algebracalliy

OpenStudy (anonymous):

f(10)=10^2+4=104 .......f(0)=0^2+4=4 ...............f(-5)=(-5)^2+4=29.........you can put in any real number, f(-100000000).f(7),f(9999999) whatever you like

OpenStudy (anonymous):

therefore the domain is any number your heart desires

OpenStudy (anonymous):

Also, it is all real numbers because there is nothing that causes this to be 0

OpenStudy (anonymous):

since x^2 function is defined for all real x ..you can put anything in place of x and calculate x^2. so there is no restriction...it is obvious so no algebraic explanation has written -infinty < x < +infinity